If I have an equation: C=500+.95(Q-T), how am I able to find C without Q? or vice versa? Thanks!
1
Expert's answer
2014-12-16T09:40:20-0500
C=500+.95(Q-T) is the equation of consumption, where Q is GDP and T is taxes. GDP = C/c, where c is marginal rate of consumption and c = 0.95 in our case.
So, to find C without Q:
C = 500 + 0.95*(C/0.95 - T) = 500 + C - 0.95T
Vice versa Q = C/0.95